OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL what is the probability of rolling a number cube and getting a 4 and then an even number?

OpenStudy (xguardians):

1/6 (Odd of getting 4) * 1/2 (Odd of even Number ) = 1/12 or 8.3%

OpenStudy (anonymous):

\[A. \frac{ 1 }{ 6 }\] \[B. \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

i didnt even put all of the answers yet lol but thanxs